OpenStudy (anonymous):

Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution

OpenStudy (anonymous):

u mean 6

hero (hero):

oops

hero (hero):

2x+8 = -6 2x+8 = 6

OpenStudy (anonymous):

^^ start zooming in hero

hero (hero):

lool

OpenStudy (anonymous):

this gives us 2 equations 2x+8=6 & 2x+8=-6 this implies x=-1,-7

hero (hero):

OS needs to invent an edit button

OpenStudy (anonymous):

Now it seems like im talking to myself lol

OpenStudy (anonymous):

@sarah u r heree.................. hii

OpenStudy (anonymous):

I concur!

hero (hero):

It's OS fault for not making an edit button

OpenStudy (anonymous):

Hey luckey =D

OpenStudy (anonymous):

@sarah hwz u dear?

OpenStudy (anonymous):

I'm good =) How are you?

OpenStudy (anonymous):

m better.............. heyy gimme ur gmail

OpenStudy (anonymous):

i don't have gmail =)

OpenStudy (anonymous):

sarah is a celebrity round here ?

hero (hero):

Apparently.

OpenStudy (anonymous):

Great!! and you are a fan ?

hero (hero):

Maybe one day I can get an autographed photo.

hero (hero):

Sorry, I had to fix grammatical errors.

OpenStudy (anonymous):

@sarah fine u dunt have any email id?

OpenStudy (anonymous):

You are forgiven Hero.

OpenStudy (anonymous):

I can't give you my email luckey =) , im only on here for math

OpenStudy (anonymous):

ok yr.............. no problem........ v can discuss the math here only